Suppose the growth of production in Country A follows Harrod-Domar Model. Country A has 100 people(labors) in Year 1. The capital K in Year 1 is 60. The incremental capital-output ratio v = 5 and the constant savings rate s = 0.1. There is no depreciation in capital. What is the growth rate of aggregate productions in Year 1?
Answer #1

In Harrod domar model

At eqm, growth rate of output

g = s/v

= .1/5

= .02

= 2%

growth rate of Aggregate Production = 2%

K = 60, L = 100

Then K/Y = 5

Y = K/5 = 60/5 = 12